OpenStudy (anonymous):

can someone explain how it goes from one step to the other? (equation posted soon)

OpenStudy (abb0t):

huh?

OpenStudy (anonymous):

\[35/85(1/6(e^(6\theta))\sin7\theta-7/36e^(6\theta)\cos(7\theta)+c\]

OpenStudy (anonymous):

then the answer from their is

OpenStudy (anonymous):

\[1/85(e^6)(6\sin(7x)-7\cos(7x))+C\]

OpenStudy (anonymous):

how does it go from the first to the second equation..what happens to the 35/38?

OpenStudy (anonymous):

*Capitol and lowercase c are the same i just happened to capitalize one and not the other sorry

OpenStudy (solomonzelman):

if you want to put up more than one symbol in the power, then do ^{ } like that.

OpenStudy (anonymous):

ohh okay thanks yea i attempted it with ^() and it was not working thanks for the info